Campus networking staff are actively investigating a problem with network traffic on one of the two campus core routers, causing intermittent connectivity problems. They are working with the vendor.
This maintenance has been completed and all services have been restored. If you notice any problems, please try rebooting the affected machines, and then email email@example.com if the issue continues after rebooting.
Our new UPS (battery backup) system is also online and working as expected (which was the purpose of this maintenance).
Power will be shut down to our primary server room in order to complete tests of and apply load to our new UPS backup system. Most services including file storage will be offline from 11PM Tues. 7/31 to 9AM Wed. 8/1.
- Network infrastructure including wired and wireless networks and DHCP/DNS services will remain online (except in 290 Soda).
- LDAP and Active Directory authentication services will remain online.
- The IRIS website and portions of the department website including personal homepages, as well as other IRIS services that depend on file storage, will be shut down starting at 11PM Tues. 7/31.
- Storage services, including home and project file storage, will be shut down at 12 AM Wed. 8/1.
- All services are scheduled to be restored by 9AM Wed. 8/1, but may be online earlier if maintenance is completed ahead of schedule.
As the IRIS website will be unavailable during this maintenance, https://status.eecs.berkeley.edu will be available for in-the-moment status updates.
Please contact firstname.lastname@example.org with any questions or concerns.
The maintenance work was completed and all services are online as of approximately 8:30 PM. Due to issues commissioning our new UPS, that task was not completed and another (shorter) shutdown will be necessary some time in the future to bring the UPS online without risk to the servers.
The EECS Network is currently disconnected from the rest of the Internet. The EECS Network team is currently investigating.
A power outage to Soda Hall which began around 12:30pm on June 21, 2017 has taken out many IRIS services. This backup website will be updated until our main website is working again.
Power has been restored to Soda Hall and staff are bringing up network equipment now. The power conditioner in the main Soda machine room is having problems recovering. Electricians are on site assisting.
This incident was fully resolved around 12:30am on June 22. See this IRIS news article for details.
Routine RedHat Linux operating system patches have been released which require system reboots. Starting at 6pm Wednesday, April 12, IDSG will be patching and rebooting affected servers. The following services will experience brief outages of 10 minutes or less between 6pm and 6:30pm.
- Department mailing list services (lists.eecs.berkeley.edu)
- Department jabber/xmpp service (jabber.eecs.berkeley.edu)
- IRIS website (iris.eecs.berkeley.edu)
The EECS networks became unreachable around 1:20pm on Monday, April 10. Staff are investigating.
Update (2:15pm): Staff investigation found the problem to be at or with one of the department border firewall devices. The secondary firewall device has been made primary now, and traffic is routing properly. Staff will further examine the problem firewall device, to narrow down the specific cause.
IRIS will be migrating our help request tracking system to a new server tonight, Monday March 27 between 6pm and 8pm. Some parts of the IRIS website will be unavailable during this time.
We ran into a number of difficulties on the electrical side which delayed restoration of services, but basic networking (wired and wireless) were brought back at approximately 12PM.
At this time we are experiencing some slowness in the LDAP system, and we have also encountered trouble in bringing the NetApp file servers online. As a result the department website and file storage, as well as other services which rely on them, are still offline.